OpenStudy (anonymous):

factor x^2-26x+169-y^2

OpenStudy (raden):

lets do step by step from x^2-26x+169-y^2 can u factor out of x^2-26x+169 (i was save the term of y^2)

OpenStudy (anonymous):

X(x-26+169-y^2)

OpenStudy (raden):

no, just factoring of x^2-26x+169, first

OpenStudy (anonymous):

1x^2-26x+169= (1x^2+13x)(13x+169)?

OpenStudy (anonymous):

1x^2-26x+169= (1x^2-13x)(13x+169)?

OpenStudy (raden):

nopes.... x^2-26x+169 = (x-13)(x-13)

OpenStudy (anonymous):

oh so (x-13)(x+13) than what do u do with the y^2

OpenStudy (raden):

it similar by (x-13)^2, right ?

OpenStudy (anonymous):

i don't know

OpenStudy (anonymous):

the answer is (x-13-y)(x+13+y)

OpenStudy (raden):

no, still incorect it should be (x-y-13)(x+y-13) sorry, my net very slow :'(

OpenStudy (anonymous):

its ok this is my new problem x^2-30x+225-y^2

OpenStudy (anonymous):

i think its (x-15-y)(x-15+y)

OpenStudy (anonymous):

its right =)

OpenStudy (anonymous):

thanks for ur help